OpenStudy (angvlinaa):

Lizzy is tiling a kitchen floor for the first time. She had a tough time at first and placed only 5 tiles the first day. She started to go faster, and by the end of day 4, she had placed 35 tiles. She worked at a steady rate after the first day. Use an equation in point-slope form to determine how many days Lizzy took to place all of the 100 tiles needed to finish the floor.

OpenStudy (thefatdog):

She placed only 5 tiles the first day and started to go faster and by theend of day 4 she had placed 35 tiles. So, we have two points: (1,5) and (4,35) Slope = 30/3 = 10 Y - 5 = m(X - 1) For 100 tiles, 100 - 35 = 10 (X - 4) 65 =10(X - 4) 6.5 = X - 4 X = 10.5 So,11th day

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\) what do u think it is